OpenStudy (anonymous):

Question 4. Which is the equation of the given line in slope-intercept form? (Points : 1)

OpenStudy (anonymous):

a. y= -1/4x + 1 b. y= -4x + 1 c. y= 4x - 1 d. y= 4x + 1

OpenStudy (danjs):

so can you calculate the slope of that line using the 2 points they have labeled for you?

OpenStudy (danjs):

given the two points on the graph (x1,y1) and (x2,y2) Calculate the slope by finding \[slope = m = \frac{ y2 - y1 }{ x2 - x1 }\]

OpenStudy (anonymous):

-1 - 1/ 5 - (-3)

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

a point is in the form (x , y) put the difference in y values on the top not bottom

OpenStudy (danjs):

(-3 - 5) / (1 -(-1) = -8 / 2 = -4

OpenStudy (danjs):

only one answer has a slope of -4

OpenStudy (anonymous):

so which one is it? @DanJS

OpenStudy (danjs):

the answers are in the form y = mx + b, where m=the slope, we figured m = -4 only one answer has that.